OpenStudy (mikey):

I will medal for answer Which of the following expressions are equivalent? Justify your reasoning. A.4√x3 B.1/ x^−1 C.10√x5•x4•x2 D.x^1/3 •x^1/3 •x^1/3

OpenStudy (mikey):

@Cas_fangirl_14

OpenStudy (cas_fangirl_14):

A. \[\sqrt[4]{ x^{3} } = x^{3/4}\] B. \[\frac{1}{ x^{-1} } = x \] C. \[ x^{1/3} * x^{1/3} * x^{1/3} = x\] D. \[x^{1/3} * x^{1/3} * x^{1/3} = x\] B and D are equivalent.

OpenStudy (mikey):

thanks so much

OpenStudy (mikey):

but I also need the justification

OpenStudy (mikey):

I need the justification so I can atleast see how u got the answer and I can understand myself

OpenStudy (mathstudent55):

The justifications for all of these are rules of exponents.

OpenStudy (mathstudent55):

\(\Large \sqrt[10]{x^5} \cdot x^4 \cdot x^2 = x^{\frac{5}{10}} \cdot x^4 \cdot x^2 =\) \(\Large =x^{\frac{1}{2}} \cdot x^4 \cdot x^2 = x^{\frac{13}{2} } = \sqrt{x^{13}} \)

OpenStudy (mathstudent55):